Question 9 : The molarity of urea in a solution prepared by dissolving 16 g of urea (MW = 60.0 g/mol) in 39 g of H2O is __________ M. The density of the solution is 1.3 g/mL.
6.3
3.7
6.8
0.11
1
Expert's answer
2018-01-31T14:16:07-0500
The total mass of solution is 39g + 16 g = 55g, and the volume of solution in liters can be calculated from the following conversion:
55 g (1mL/1.3g)(1L/1000mL) = 0.042 L.
The molarity (C) of solution is given by a formula:
C = m/(MV),
where m is the mass solute, M is the molar mass of solute and V is the volume of solution.
C = 16g/(60.0 g/mol · 0.042 L) = 6.3 mol/L.